Solve the system of equations by using the addition method:
4x + 3y = - 5
-4x - 10y = 26
Please show steps. Thank you.

Answer by Alan3354(69443)

Solve the system of equations by using the addition method:
4x + 3y = - 5
-4x - 10y = 26
------------------ Add them